OpenStudy (anonymous):

Need some help ASAP with some math. I'll post what I have done so far and I'll post the problem in the comments. What you know is: Percentage of A: 60% (60/100 as a fraction) Percentage of B: 85% (85/100 as a fraction) Final percentage: 80% (80/100 as a fraction) Total Milliliters: 150 mL (Percentage of dressing A) x (milliliters you need) + (Percentage of dressing B) x (milliliters you need) = (final percentage) x (Total milliliters) Now you actually plug things in (60/100) x A + (85/100) x B = (80/100) x 150 To solve: First combine the things that do not involve a variable

OpenStudy (anonymous):

The problem

OpenStudy (anonymous):

To solve: First combine the things that do not involve a variable (80/100) x 150 80x150 =12,000 Divided by 100 = 120 So now the equation reads (60/100) x A + (85/100) x B = 120

OpenStudy (anonymous):

Now you are not able to solve for two variables at one time, so you need to eliminate one. To do this you use substitution. Since you know that A+B= 150 Meaning, you will use A mL of dressing A and B mL of dressing B to get a total of 150 mL So you get A by itself, B= 150-A Now plug this into your equation where B was (60/100) x A + (85/100) x (150-A) = 120 Now you can distribute: (60/100)A + (80/100 x 150) – (85/100 x A) = 120

OpenStudy (anonymous):

And then I'm lost on what to do next

OpenStudy (mertsj):

Amount of 60% solution to use: x Amount of 85% solution to use: 150-x .6x+.85(150-x)=.80(150)

OpenStudy (mertsj):

.6x+127.5-,85x=120 -.25x=-7.5 x=30 pints of 60 % solution 150-30=120 pints of 85% solution

OpenStudy (anonymous):

I'm sorry, I'm still totally lost. :( Math learning disabilities suck.

OpenStudy (anonymous):

Somebody please help me out.

OpenStudy (mertsj):

If you have 5 pounds of candy and 3 pounds of it is snickers and the rest is jelly beans wouldn't the jelly beans be 5-3?